OpenStudy (anonymous):

(6m)^5 x (6m)^6 Multiply and simplify

OpenStudy (cwrw238):

= 6^11* m^11 = 362797056m^11

OpenStudy (mertsj):

Remember, when multiplying, add the exponents: \[(6m)^5(6m)^6=(6m)^{6+5}=(6m)^{11}\]